Find the critical values chi squared Subscript Upper L χ2L and chi squared Subscript Upper R χ2R for the given confidence level c and sample size n. c= 0.95, n= 26
Answer #1

We have here, n = 26 and c= 0.95

Degree of freedom = n - 1 = 26- 1 = 25

......................by using Chi square critical value calculator or Excel command =CHIINV(1-(0.05/2),25)

..............by using Chi square critical value calculator or Excel command =CHIINV((0.05/2),25)